a rectangular box is to be used to store materials. The inside of the box is 12.0 in. long, 9.50 in. wide, and 8.75 in. deep. What is the area of sheet lead that must be used to line the inside of the box?
Answer by checkley71(8403)

BOTT0M MEASUREMENT=12*9.5=114 SQ INCHES.
2 OF THE SIDES=2*12*8.75=210 SQ INCHES.
THE OTHER 2 SIDES=2*9.5*8.75=166.25 SQ INCHES.
114+210+166.25=490.25 SQ INCHES REQUIRED TO LINE THIS BOX..